Thomas has a set of 26 alphabet cards, each having a different letter of the alphabet on it. He randomly draws one card and then rolls a fair die, numbered from 1 to 6. What is the probability he draws a card with a vowel (a, e, i, o, u) on it and then rolls an odd number?


Answer by sudhanshu_kmr(1152)


probability of draw a card with a vowel = 5/26 ( any one of 5 vowels)

probability of getting an odd number = 3/6 = 1/2 (any one of 3 odd number)

so, required probability = 5/26 * 1/2 = 5/52